OpenStudy (anonymous):

Okay please help :3 Find the 12th term of the sequence which has a first term of 14 and has a common ratio of 1.03. Round your answer to the nearest hundredth if necessary.

OpenStudy (raden):

use the formula : Tn = a1 * r^(n-1) in this case, n = 12, r = 1.03, also given a1 = 14 now plugin all values to formula above

OpenStudy (anonymous):

I got 5.6055452 E12 uhhh that doesn't even sound right.

OpenStudy (anonymous):

*5.60554527

OpenStudy (raden):

i got different like yours

OpenStudy (raden):

can you recheck again your calculation Tn = a1 * r^(n-1) T12 = a1 * r^(12-1) T12 = 14 * (1.03)^11 T12 = 14 * 1.384 T12 = ..... ?

OpenStudy (anonymous):

Well following your steps the answer would be 19.376

OpenStudy (anonymous):

Oh I think I know where I went wrong. I didn't do (1.03)^11 separately

OpenStudy (raden):

yup... do you have the answer choices ?

OpenStudy (anonymous):

It is not multiple choice question.

OpenStudy (anonymous):

:(

OpenStudy (raden):

well, must be sure with the answer :v

OpenStudy (anonymous):

I got answer correct! Thank you very much! :D

OpenStudy (raden):

you're welcome :*